Michael is receiving an annuity due with monthly payments for 20 years. Each monthly payment in the first year is 130. Each monthly payment in the second year is 260. Each monthly payment in the third year is 390. The payments continue to increase in the same pattern until each monthly payment in the 20th year is 2600. Using an annual effective rate of interest of 7%, calculate the present value of this annuity.
Answer #1

02) C+O7) -I. O.O0 5654145 12 PV= P 20 130C a201 12 20 V 12 12 12 -20 C1-01)07)= \1-33559524 1 O.O7 20 - 20 20V = 20(1-07) 15